if f(x)=a raise to x,a>0,show that
f(x+1)=a*f(x)=f(2)*f(x-1)
1
Expert's answer
2014-05-01T04:58:19-0400
We have
f(x+1)=ax+1=a*ax=a*f(x) - this is the fist equality
For the second equality we have
f(2)*f(x-1)=a2ax-1=a*(a*ax-1)=a*ax=a*f(x).
QED.